OpenStudy (anonymous):

Can ya'll help me factor this expression completely? x^2 -9

OpenStudy (debbieg):

I assume you're supposed to factor it? This is a special product, you might (hopefully :) remember from when you learned to FOIL/do binomial multiplication, that: \(\Large (a-b)(a+b)=a^2-b^2\) E.g., the product of a sum and a difference is = to the difference of the squares of the first and last terms. So reverse that to factor: If you have a difference of squares, it factors into the product of the sum and the difference of the square roots: \(\Large a^2-b^2=(a-b)(a+b)\)